# Think in Safety: Unveiling and Mitigating Safety Alignment Collapse in Multimodal Large Reasoning Model

### v1

@@ C.1 Data Statistics
C.2 Prompt for Data Conustruction
@@ D.1 Detailed Experiments on MSSBench
In the evaluation protocol of MSS-Bench, scenarios are divided into chat and embodied settings, with safety scores calculated separately for safe and unsafe conditions. The average safety scores are reported in Table 2, and the specific scores for each category are detailed in Table 7.
@@
D.2 Case Study
Examples of responses generated by fine-tuned on VLGuard, MIS, and our dataset are illustrated in Figure 7. Our approach demonstrates the ability to retain the thought process of the models while decisively rejecting unsafe inputs and explicitly articulating the potential serious consequences associated with such queries.
Figure 7: Examples of several baselines and our dataset on $M$-SafetyBench. We only highlight some fragments to show the representative difference between various methods. The red indicates the unsafe parts, while the blue indicates the safe parts.

### v2

@@ B.2 Dataset Statics
B.3 Prompt for Data Construction
@@
Appendix C Training Details
C.1 Baselines
To evaluate the effectiveness of our data construction, we select three existing multimodal safety alignment datasets as baselines that are listed as follows:.
  • SPA-VL Zhang et al. (2024): A safety preference alignment dataset for Vision Language Models.
C.2 Traning Details
For the training procedure, we utilize the LoRA Hu et al. (2022) to fine-tune the R1-Onevision and LLaVA-CoT based on the LLaMA-Factory [^1] https://github.com/hiyouga/LLaMA-Factory framework. For SPA-VL dataset, we use the DPO script for training. All training experiments are conducted using one NVIDIA A100-80G GPUs. The hyper-parameters for supervised fine-tuning are listed in Table 9.
Hyper-Parameter Value
lora rank 8
learning rate 1.0e-5
train epoch 1.5
per_device_batchsize 1
warm up ratio 0.1
learning rate scheduler cosine
Table 9: Hyper-parameters for supervised fine-tuning.
@@ D.1 Detailed Experiments on MSSBench
In the evaluation protocol of MSSBench, scenarios are divided into chat and embodied settings, with safety scores calculated separately for safe and unsafe conditions. The average safety scores are reported in Table 1, and the specific scores for each category are detailed in Table 7.
D.2 Performance of Thought Process on MM-SafetyBench
We also independently evaluate the safety performance of the thought process on MM-SafetyBench. This section presents the MLRMs’ performance and fine-tuned models’ performance in Table 10 and Table 11.
  R1-Onevision Safe Answer Unsafe Answer
Safe Thought 12.65% 1.90%
Unsafe Thought 7.34% 78.11%
Kimi-VL-Thinking Safe Answer Unsafe Answer
Safe Thought 12.68% 2.74%
Unsafe Thought 25.83% 58.75%
LLaVA-CoT Safe Answer Unsafe Answer
Safe Thought 15.89% 3.06%
Unsafe Thought 11.85% 69.23%
QVQ-Preview Safe Answer Unsafe Answer
Safe Thought 24.52% 6.91%
Unsafe Thought 6.19% 62.38%
Skywork-R1V Safe Answer Unsafe Answer
Safe Thought 19.88% 3.93%
Unsafe Thought 7.44% 68.75%
Table 10: Proportion of safety in the thought process and the final answer on MM-SafetyBench.
  R1-Onevision Safe Answer Unsafe Answer
Safe Thought 12.65% 1.90%
Unsafe Thought 7.34% 78.11%
R1-Onevision+Ours Safe Answer Unsafe Answer
Safe Thought 74.58% 8.87%
Unsafe Thought 16.55% 0.00%
LLaVA-CoT Safe Answer Unsafe Answer
Safe Thought 15.89% 3.06%
Unsafe Thought 11.85% 69.23%
LLaVA-CoT+Ours Safe Answer Unsafe Answer
Safe Thought 71.73% 21.73%
Unsafe Thought 6.43% 0.06%
Table 11: Proportion of safety in the thought process and the final answer of fine-tuned model using TiS on MM-SafetyBench.
@@ D.2 Performance of Thought Process on MM-SafetyBench
D.3 Case Study of Ablation Experiments
Examples of responses generated by fine-tuned on TiS and TiS no thought dataset are illustrated in Figure 8 and Figure 9. The TiS dataset without thought is predominantly composed of unsafe examples. When fine-tuning MLRMs based on this data, the model tends to generate responses that consist solely of brief refusals. In particular, for safety-awareness examples, the model tends to either produce overly brief refusals or fail to detect unsafe risks. In contrary, the complete TiS data enables the detection of more unsafe intents during the thought process, while also providing more helpful information.
Figure 8: Examples from R1-Onevision tuning on TiS w/o thought process. The red indicates the unsafe parts, while the blue indicates the safe parts.
Figure 9: Examples from LLaVA-CoT tuning on TiS w/o thought process. The red indicates the unsafe parts, while the blue indicates the safe parts.
D.4 Case Study
Examples of responses generated by fine-tuned on VLGuard, MIS, SPA-VL and TiS dataset are illustrated in Figure 10. Our approach demonstrates the ability to retain the thought process of the models while decisively rejecting unsafe inputs and explicitly articulating the potential serious consequences associated with such queries.
Figure 10: Examples of several baselines and our dataset on MM-SafetyBench. We only highlight some fragments to show the representative difference between various methods. The red indicates the unsafe parts, while the blue indicates the safe parts.
Figure 11: Examples of several baselines and our dataset on SIUO. We only highlight some fragments to show the representative difference between various methods. The red indicates the unsafe parts, while the blue indicates the safe parts.